# Hybrid Deepfake Image Detection: A Comprehensive Dataset-Driven Approach Integrating Convolutional and Attention Mechanisms with Frequency Domain Features

### v1

@@ 2.4.2 Data Efficient Image Transformer (DeiT) Model
Figure 6: Data Efficient Image Transformer (DeiT) architecture
Note that CNN models primarily capture local contexts and eventually may capture global contexts through hierarchical learning. To this end, DeiT is an effective alternative to traditional CNNs in learning global features [6]. It uses knowledge distillation techniques and optimization strategies to improve generalization. DeiT learns from a pre-trained CNN model (that is, the RegNetY-16GF model [39]) and uses a distillation token during training, allowing it to learn faster and more efficiently with less data [36].
@@
Here $c$ is the class index. Using this prediction, the final loss function for hard distillation is treated as
@@
The first term of the equation can be thought of as $L_{\text{cross-entropy}}$ and the second term as $L_{\text{teacher}}$ depicted in Figure 6
As the competition dataset lacks enough samples to generalize a ViT, we chose DeiT to capture global contexts effectively, despite it being computationally costly. There exist different DeiT model variations based on the number of heads and embedding resolutions. In our proposed solution, we have implemented the DeiT-B model, which uses the ViT-B model as its backbone. It has 768 embedding dimensions with 12 heads containing a total of 86 million parameters [36]. The DeiT model architecture is shown in Figure 6 and the number of parameters and the output shapes of each layer are shown in Table 2.
Table 2: Details of the Data Efficient Image Transformer (DeiT) Model
@@
Linear-2 $\times$ 12 [197, 768] 35,407,872
Attention $\times$ 12 [197, 768] 0
Linear-3 $\times$ 12 [197, 3072] 28,348,416
GELU $\times$ 12 [197, 3072] 0
@@
2.4.3 XceptionNet Model
XceptionNet is based on the concept of depthwise separable convolutions, which enhance computational efficiency while maintaining accuracy. In depthwise convolution, a single convolutional filter is applied to each input channel. Following this, pointwise convolution is used to combine the outputs of the depthwise convolution into a linear form [40]. XceptionNet features a fully convolutional architecture for simplicity and efficiency, leverages residual connections to improve gradient flow, and employs deep feature extraction.
(a) Entry Flow
(b) Middle Flow
(c) Exit Flow
Figure 7: Details of XceptionNet architecture
As shown in Figure 7, our implemented XceptionNet has 36 convolutional layers, structured into 14 modules with residual connections, except for the first and last modules [40]. It has three primary sections—Entry Flow, Middle Flow, and Exit Flow. The Entry Flow layer consists of convolutional layers followed by max pooling and is responsible for reducing spatial dimensions and extracting low-level features. The Middle Flow layer contains multiple depthwise convolution blocks with residual connection, focusing on high-level features. The Exit Flow layer includes more depthwise separable convolutions and reduces spatial dimensions further. We modified the fully connected layer from the original architecture to match our binary classification problem. The summary of the output shapes and number of parameters of each block and layer is shown in Table 3.
Table 3: Details of the XceptionNet Model
Layer (type) Output Shape Param #
Conv2d [32, 149, 149] 864
BatchNorm2d + ReLU [32, 149, 149] 64
Conv2d [64, 147, 147] 18,432
BatchNorm2d + ReLU [64, 147, 147] 128
Block-1 [128, 74, 74] 35,264
   —SeparableConv2d [128, 147, 147] 8,768
   —BatchNorm2d + ReLU [128, 147, 147] 256
   —SeparableConv2d [128, 147, 147] 17,536
   —BatchNorm2d [128, 147, 147] 256
   —MaxPool2d [128, 74, 74] 0
   —Conv2d [128, 74, 74] 8,192
   —BatchNorm2d [128, 74, 74] 256
Block-2 [256, 37, 37] 136,064
Block-3 [728, 19, 19] 915,944
Block-4 $\times$ 8 [728, 19, 19] 12,911,808
Block-5 [1024, 10, 10] 2,039.584
SeparableConv2d [1536, 10, 10] 1,582,080
BatchNorm2d + ReLU [1536, 10, 10] 3072
SeparableConv2d [2048, 10, 10] 3,159,552
BatchNorm2d + ReLU [2048, 10, 10] 4096
SelectAdaptivePool2d [2048] 0
Linear [2] 4098
Total Parameters 20,811,050
Trainable Parameters 20,811,050
@@ 2.4.3 XceptionNet Model
2.4.4 Weighted Ensemble Mechanism
Finally, we implemented a weighted ensemble mechanism to combine the predictions of the aforementioned three models. The associated weights for the ResNet-34 with SE, DeiT, and Xception with wavelet transform are 0.36, 0.45, and 0.19, respectively, based on our parameter search. The final architecture is shown in Figure 8, and the total number of trainable and non-trainable parameters in the final architecture is shown in Table 4.
Figure 8: The proposed weighted ensemble mechanism
Table 4: Number of parameters in the final weighted ensemble mechanism
Model Total Parameters Trainable Non-trainable
ResNet 21,319,010 8,204,642 13,114,368
DeiT 85,648,130 85,648,130 0
Xception 20,811,050 20,811,050 0
Total 127,778,990 114,664,622 13,114,368
@@ 3 Experimental Results
Figure 9: (a) Confusion matrix and (b) ROC curve for validation set
In this section, we demonstrate the performance of our proposed solution on the competition dataset. We note that investigation and experimentation with different architectures helped us make informed decisions to reach our proposed solution. More specifically, we started experimenting with simpler CNN models. As mentioned before, the performance of CNN models, such as the VGG-19, indicated insufficient capacity and indicated the necessity of deeper networks. Additionally, residual networks, such as the ResNet-50, raised overfitting issues. These results motivated us to choose a ResNet-34-based architecture to strike a balance between high-bias and high-variance regimes. We investigated the performance of the convolutional block attention module (CBAM) and SE and opted for SE block based on the generalization performance on the given dataset. Lastly, we investigated the effect of different handcrafted feature extraction methods on the performance and opted for wavelet feature extraction for our XceptionNet model.
(a) SEResNet (Untrained)
(b) DeiT (Untrained)
(c) Xception (Untrained)
(d) SEResNet (Trained)
(e) DeiT (Trained)
(f) Xception (Trained)
In essence, gradually increasing the complexity of networks helped us understand the shortcomings of different architectures for our specified task and led us to choose the model architectures based on their generalization capability with a relatively small dataset. To further reduce the variance, we employed a weighted ensemble of ResNet-34 with SE, DeiT, and XceptionNet to generate the final results. We present a comparative analysis of the results of different model variations and configurations that guided us to our final design decisions in Table. 5. We argue that these results serve as empirical evidence of our choice of neural network architectures, and the weighted ensemble approach.
@@
Model Validation Acc.
Xception-Net [40] 61.48%
@@
ResNet-34 [4] 80.32%
ResNet-34 with CBAM 85.19%
Xception-Net with Wavelet Transform 87.76%
ResNet-34 with SE 88.90%
@@
Maj. Voting (ResNet, DeiT, Xception) 91.34%
Weighted Ensem. (ResNet, DeiT) 91.80%
Weighted Ensem. (ResNet, DeiT & Xception) (Proposed) 93.23%
Hyperparameters. For the training of our models, we have used the ADAM optimizer. We have used accuracy, precision, recall, and F1 score as our evaluation metrics. Since the deepfake detection problem is a binary classification problem, we used binary categorical cross-entropy as our loss function. Using the validation set to observe the accuracy of the proposed architecture, each image took approximately 0.03515625 sec on a Tesla P100 GPU. The hyperparameters used for training are presented in Table 6.
@@
Number of Epochs
   - ResNet-34 with SE 10
@@
We have utilized the NumPy and Pandas libraries for preprocessing and data manipulation tasks, and PyTorch [42] for model implementation, training, and validation. More specifically, we used torch for core functionality, torch.nn for neural networks, torch.optim for optimization, and utilities like torch.utils.data.Dataset, DataLoader, and random_split. From TorchVision, we used transforms for image transformations and models for pre-trained architectures. For performance metrics, we applied Scikit-learn (precision_score, recall_score, and f1_score). Progress tracking was managed with TQDM, and TorchSummary provided model summaries. Image processing utilities include OpenCV, while PyWavelets was utilized for wavelet transformations. Additionally, Timm was used to access advanced pre-trained models, and the Random module was applied for generating randomness.
(a) SEResNet
@@
Result Analysis. As mentioned before, we evaluated our weighted ensemble model for multiple evaluation metrics, including accuracy, precision, recall, F1-score, area under receiver operating characteristic (AU-ROC) curve, equal error rate, and average precision. While calculating these metrics, fake images were considered as class 0 and real images as class 1. The performance of our model on these metrics is presented in Table 7 and the confusion matrix and the ROC curve are shown in Figure 9. It is evident from the evaluation indices that our proposed weighted ensemble architecture performs very well on the competition dataset, which underlines the generalization capability of the proposed approach.
Table 7: Performance of the Final Ensemble Architecture combining ResNet-34, DeiT, and XceptionNet
Evaluation Metrics Performance
Accuracy 93.23%
Precision 0.9515
Recall 0.9121
F1 score 0.9314
Area under ROC 0.9769
Equal Error Rate 6.33%
Average Precision 0.9744
To investigate the performance of our ensemble model, we intend to visualize i) the regions of the images that the three models focus for classification, and ii) how the three models distinguish real and fake images before and after training. To that end, we note that the Grad-CAM visualizations show how different models focus on different facial regions of the same image for deepfake detection. In Figure 11, it can be seen that the ResNet-34 and Xception models focus on localized facial features, whereas DeiT displays diffuse attention patterns indicating a focus on global features. Each of the models focuses on different parts of the image, capturing unique and complementary information. When these diverse insights are combined into an ensemble, the overall prediction becomes more accurate, robust, and better at generalizing to new data. Additionally, we present the Grad-CAM view of some of the wrong predictions by the three models in Figure 12, showing the focus of irrelevant regions, leading to mis-classifications.
@@
Next, we note that high-dimensional data can be visualized by projecting onto a two-dimensional space using the non-linear dimensionality reduction technique known as t-SNE (t-Distributed Stochastic Neighbor Embedding) [43]. By maintaining the local neighborhoods of data points, it shows the differences and similarities of data points in the learned feature space. As such, the t-SNE visualizations (Fig. 10) offer valuable insights into the effectiveness of our models. For both real and fake images, we observe a single cohesive cluster for all three models. Notably, the cluster of fake images generated by two of our models, SEResNet and XceptionNet, is compact, despite these images being produced using various deepfake generation methods. The plots show that the training has effectively enhanced the ability to distinguish between real and fake image features for all three models. This, in-turn, highlights the generalization capability of our proposed ensemble model. After training, all three backbone models—SEResNet, DeiT, and XceptionNet—produced well-separated embeddings for real and fake images, as illustrated in Fig. 10(d), 10(e), and 10(f).

### v2

@@ 2.4.2 Data Efficient Image Transformer (DeiT)
DeiT is an effective alternative to traditional CNNs in learning global features [22], using knowledge distillation techniques and optimization strategies to improve generalization. DeiT learns from a pre-trained CNN model (that is, the RegNetY-16GF model [42]) and uses a distillation token during training, allowing it to learn faster and more efficiently with less data [39].
@@
Here, $c$ is the class index. Using this prediction, the final loss function for hard distillation is treated as
@@
The first term of the equation can be thought of as $L_{\text{cross-entropy}}$ and the second term as $L_{\text{teacher}}$.
Our proposed solution utilizes the DeiT-B model, a 768-dimensional model with 12 heads and 86 million parameters, based on the ViT-B model [39]. The summary number of parameters and the output shapes of layers are shown in Table 2.
Table 2: Summary of the Data Efficient Image Transformer (DeiT) Model Architecture
@@
(Linear-2 + Attention) $\times$ 12 [197, 768] 35,407,872
(Linear-3 + GELU) $\times$ 12 [197, 3072] 28,348,416
@@
2.4.3 ConvNeXt Model
ConvNeXt combines the efficiency and inductive biases of traditional CNNs with the scalability and performance of transformers. It employs an inverted bottleneck block, similar to transformers’ MLP blocks, reducing FLOPs in downsampling layers. Here, batch normalization is replaced with layer normalization, aligning with transformer practices, and GELU is used instead of ReLU for smoother non-linearity [43].
Inspired by both ResNets [44] and Swin transformers [45], the architecture includes a patchify stem where input images are processed by a 4x4 convolution with stride 4, followed by layer normalization, to downsample the image into non-overlapping patches. This mimics ViT’s patch embedding but uses convolutions.
Each ConvNeXt block contains a 7x7 depthwise convolution (one filter per input channel) that is applied to capture spatial features efficiently. Then layer normalization is applied before and after the convolution to stabilize training. It normalizes features across the channel dimension for each spatial position. For feature $X[i,j,:]~{}\epsilon~{}\mathbf{R^{C}}$ the normalized output is:
$\hat{X}[i,j,c]=\frac{X[i,j,c]-\mu}{\sqrt{\sigma^{2}+\epsilon}}\cdot\gamma+\beta$ (11)
Here $\mu=\frac{1}{C}\sum_{c=1}^{C}X[i,j,c]$, $\sigma^{2}=\frac{1}{C}\sum_{c=1}^{C}(X[i,j,c]-\mu)^{2}$ , $\epsilon$ is a small constant, and $\gamma$ and $\beta$ are trainable parameters. This type of normalization works better than typical batch normalization [46]. After that, pointwise convolution is applied to expand (to 4x the input dimension) and contract the feature map, mimicking transformer MLP layers. GELU (Gaussian Error Linear Unit) [47] is applied instead of RELU for smoother non-linearity.
$\text{GELU}(x)=x\cdot\Phi(x)=x\cdot\frac{1}{2}[1+\text{erf}(\frac{x}{\sqrt{2}})]$ (12)
A skip connection is used from the input to the block’s output, similar to ResNet’s residual connections [44]. Between stages, ConvNeXt uses a downsampling layer with layer normalization and a 2x2 convolution with stride 2, inspired by Swin Transformer’s separate downsampling strategy, unlike ResNet’s within-block downsampling. From the diverse family of ConvNeXt architectures, we specifically implemented the ConvNeXt-tiny model for its low compute and high accuracy on the dataset. A summary of the different layers in the ConvNeXt-tiny architecture that we have implemented is presented in Table 3.
Table 3: Summary of the ConvNeXt Architecture
Layer Type Output Shape Param #
Conv2d + LayerNorm2d $[96,74,74]$ 4,896
CNBlock (×$\times 3$) $[96,74,74]$ 237,600
LayerNorm2d + Conv2d $[192,37,37]$ 74,112
CNBlock (×$\times 3$) $[192,37,37]$ 917,568
LayerNorm2d + Conv2d $[384,18,18]$ 295,680
CNBlock (×$\times 9$) $[384,18,18]$ 10,813,824
LayerNorm2d + Conv2d $[768,9,9]$ 1,181,184
CNBlock (×$\times 3$) $[768,9,9]$ 14,287,104
AdaptiveAvgPool2d + LayerNorm2d $[768,1,1]$ 1536
Flatten + Linear $[1]$ 769
Total Parameters 27,814,273
Trainable Parameters 27,814,273
@@ 2.4.3 ConvNeXt Model
2.4.4 Proposed Weighted Ensemble Architecture: CAE-Net
Finally, we implemented a weighted ensemble mechanism to combine the predictions of the aforementioned three models. The associated weights for the EfficientNet, DeiT, and ConvNeXt with wavelet transform are 0.35, 0.34, and 0.31, respectively, based on our parameter search. The final architecture is shown in Figure 3.
Figure 3: The proposed CAE-Net architecture
2.5 Adversarial Perturbations
Adversarial perturbations are crafted noises or modifications added to classifier model inputs, often imperceptible to humans, that can significantly deteriorate model performance and lead to misclassification. Such adversarial perturbations can be universal or image-dependent [48] and are usually done in white-box settings where access to the model’s architecture and parameters is required. Also, adversarial attacks on a known network have been shown to work effectively against other networks as well [49], making black-box attacks an imminent threat for sensitive issues like deepfake detection [50]. We evaluate the robustness of our proposed models by testing them against a simple but effective adversarial perturbation, the Fast Sign Gradient Method (FGSM). FGSM is a white-box attack without specifying a target class. Using this method, adversarial images are created by calculating the gradient of loss for input data and adjusting the input by a small amount in the direction of the gradient sign [49].
$x_{\text{adv}}=x+\epsilon\cdot\text{sign}(\nabla_{x}J(\theta,x,y))$ (13)
Here $(x,y)$ are the original input image and its label, $\theta$ indicates model parameters, $x_{adv}$ is the adversarially perturbed image, and $\nabla_{x}J(\theta,x,y)$ is the gradient of the loss function $J$ with respect to the input image $x$. The parameter $\epsilon$ controls the magnitude of the perturbation added to the original image. This method works like a gradient ascent algorithm trying to maximize the loss with the smallest possible perturbation, $\epsilon$, which makes it barely noticeable to the human eye. Though FGSM is one of the simplest forms of perturbations, it has been shown to reduce the accuracy of high-quality detectors to a large extent [51]. Its effectiveness for binary classification problems makes it an ideal choice for testing the robustness of our proposed method.
@@ 3 Experimental Results
In this section, we demonstrate the performance of our proposed solution on the competition dataset. No prior work exists on this novel combined deepfake image detection dataset, which consists of images from eight different benchmark deepfake video datasets. To address this, we hypothesized that lightweight CNNs(e.g., EfficientNet) would excel in resource-constrained settings, modern CNNs (e.g., ConvNeXt) would leverage transformer-inspired designs for better feature extraction, and ViTs would capture the global context of images. We systematically evaluated these models to test these hypotheses along with multiple other frequency domain feature extraction techniques, establishing the first performance benchmarks for the dataset.
@@
Classifier Models Accuracy
CNN models
@@
  MobileNet-V2 78.12%
  ResNet-34 [44] 80.32%
  RegNet [53] 80.92%
  Inception V3 [54] 84.38%
  ResNet-34 with CBAM 85.19%
  ResNet-34 with single SE 88.90%
  DenseNet-121 [55] 90.01%
  EfficientNet-B0 [21] 90.79%
  EfficientNet-B4 [21] 87.27%
  EfficientNet-B5 [21] 88.31%
  ResNeXt [56] 85.19%
  ConvNeXt [43] 88.18%
Frequency Domain Experimentations
@@
  ResNet-34 with Wavelet 80.44%
  EfficientNetB0 with Wavelet 82.48%
  Xception-Net [57] with Wavelet 87.76%
  ConvNeXt with Wavelet 89.49%
Study on Transformers
  ViT-B [58] 78.12%
  Swin Transformer [45] 86.52%
  MobileViT [59] 87.04%
@@
Ensemble Results
  Maj. Voting (EfficientNet, DeiT, ConvNeXt) 94.24%
  Proposed Weighted Ensem. (EfficientNet, DeiT, ConvNeXt) 94.63%
Hyperparameters. For the training of our models, we have used the ADAM optimizer and binary categorical cross-entropy as our loss function. Using the validation set to observe the accuracy of the proposed architecture, each image took approximately 0.03385416s on a Tesla P100 GPU. The hyperparameters used for training are presented in Table 5.
@@
Number of Epochs in each session
   - EfficientNet 5
@@
Result Analysis. We evaluated our weighted ensemble model for multiple evaluation metrics, including accuracy, precision, recall, F1-score, area under the receiver operating characteristic (AU-ROC) curve, equal error rate, and average precision. While calculating these metrics, real images were considered as class 0 and fake images as class 1. The performance of our model on these metrics is presented in Table 6 and the confusion matrix and the ROC curve are shown in Figure 4. It is evident from the evaluation indices that our proposed weighted ensemble architecture performs very well on the competition dataset, which underlines the generalization capability of the proposed approach.
Table 6: Performance of the Final Ensemble Architecture combining ConvNext, DeiT, and EfficientNet
Evaluation Metrics Performance
Accuracy 94.63%
Precision 0.937
Recall 0.956
F1 score 0.9464
Area under ROC 0.9737
Equal Error Rate 4.72%
Average Precision 0.9771
Figure 4: (a) Confusion matrix and (b) ROC curve for validation set
To investigate the performance of our ensemble model, we intend to visualize i) the regions of the images that the three models focus on for classification, and ii) how the three models distinguish real and fake images before and after training. To that end, we note that the Grad-CAM visualizations show how different models focus on different facial regions of the same image for deepfake detection. In Figure 5, it can be seen that EfficientNet focuses on localized facial features, whereas DeiT displays diffuse attention patterns indicating a focus on global features. When the diverse insights from each model are combined into an ensemble, the overall prediction becomes more accurate, robust, and better at generalizing to new data. Additionally, we present the Grad-CAM view of some of the wrong predictions by the three models in Figure 6, showing the focus of irrelevant regions, leading to misclassifications.
(a) EfficientNet
@@
(a) EfficientNet (Untrained)
(b) DeiT (Untrained)
(c) ConvNeXt (Untrained)
(d) EfficientNet (Trained)
(e) DeiT (Trained)
(f) ConvNeXt (Trained)
@@
Next, we note that high-dimensional data can be visualized by projecting onto a two-dimensional space using the non-linear dimensionality reduction technique known as t-SNE (t-Distributed Stochastic Neighbor Embedding) [60]. By maintaining the local neighborhoods of data points, it shows the differences and similarities of data points in the learned feature space. As such, the t-SNE visualizations (Fig. 7) offer valuable insights into the effectiveness of our models. For both real and fake images, we observe cohesive clusters for all three models, despite the images being produced using various deepfake generation methods. This highlights the generalization capability of our proposed ensemble model. After training, all three backbone models, i.e., EfficientNet, DeiT, and ConvNeXt, produced well-separated embeddings for real and fake images, as illustrated in Fig. 7(d), 7(e), and 7(f). There exist very few works on deepfake detection that perform well on the combination of such a diverse set of datasets. A comparative analysis among existing works with our proposed architecture is shown in Table 7. Note that the AUC score presented here is the mean score of these models on the corresponding datasets (we refer the reader to Figure 8 of [25]). We want to emphasize that existing works often demonstrate good performance on only a few datasets, but the overall performance on multiple datasets is very poor for most architectures.
Table 7: Comparative Analysis of Average AUC Score of Different Models
[…]
DFD [27], DFDC [31], Celeb-DF [25]
Meso4 [63] UADFV [29], DF-TIMIT [62], FaceForensics++ [26], 75.9% 10.63%
DFD [27], DFDC [31], Celeb-DF [25]
FWA [64] UADFV [29], DF-TIMIT [62], FaceForensics++ [26], 82.1% 14.45%
DFD [27], DFDC [31], Celeb-DF [25]
Xception-C23 [26] UADFV [29], DF-TIMIT [62], FaceForensics++ [26], 86.4% 11.96%
DFD [27], DFDC [31], Celeb-DF [25]
DSP-FWA [25] UADFV [29], DF-TIMIT [62], FaceForensics++ [26], 87.4% 12.79%
DFD [27], DFDC [31], Celeb-DF [25]
Our Proposed Method Celeb-DF [25], DFDC [31], DFD [27], UADFV [29], 97.37% -
FaceShifter [28], DFDC Preview [30], FaceForensics++ [26]
Finally, we look into the effectiveness of our ensemble method through a simple adversarial perturbation attack, such as Fast Gradien Sign Method (FGSM). As mentioned before, FGSM can fool models by adding simple perturbation noise based on the gradient sign that intends to maximize the loss function of the detection model instead of minimizing it. We present the accuracy scores of the separate detection model as well as the ensemble models for different values of $\epsilon$ in Table 8.
Table 8: Comparison of Our Proposed Architectures under Adversarial FGSM Attacks
Model Accuracy
$\epsilon=0.01$ $\epsilon=0.03$ $\epsilon=0.05$
EfficientNetB0 38.54% 46.48% 53.94%
Data Efficient Image Transformer 42.76% 21.94% 16.73%
ConvNeXt with wavelet 6.45% 6.51% 15.40%
Weighted Ensem.(EfficientNet, DeiT, ConvNeXt) 12.04% 13.48% 16.05%
Majority Voting (EfficientNet, DeiT, ConvNeXt) 25.52% 20.41% 19.76%
The results show an increased robustness of the EfficientNet architecture, which depicts increased accuracy with increasing perturbation compared to the other two architectures. DeiT shows a decrease in accuracy with rising perturbation, and ConvNeXt accuracy drops significantly even for very low perturbation added to the image. Though a weighted ensemble model is appropriate for achieving high accuracy on the generalized dataset, we argue that it is highly vulnerable to adversarial attacks. If any model performs poorly on adversarial data, then the performance of the entire ensemble drops significantly. However, in the majority voting ensemble, the performance does not reduce as much, since it requires the majority of models to perform poorly, implying increased robustness against adversarial attacks for applications, such as deepfake detection.